Find the limiting reagent if you have 23 moles of C, 18 miles of B and 13.2 moles of A if you have 3A+4B+5C=6ABC


1
Expert's answer
2021-05-14T02:34:55-0400

In this reaction we have to use the method of limiting reagent .

So first of all in the reactant side check which one is limiting reagent.To obtain the limiting reagent , divide the given moles of the reactants with its respective stoichiometric coffecient. The reactant which gives the least value is the limiting reagent.

In the given case -

A has 13.2 moles and its stoichiometric coffecient is 3 , so, 13.2÷3=4.4

B has 18 moles and its stoichiometric coff. is 4, so, 18÷4=4.5

C has 23 moles and its stoichiometric coff. is 5, so, 23÷5=4.6

As C gives the least value , C is the limiting reagent.

Hence the limiting reagent is C